OpenStudy (bossimbacon):

What is the solution to the equation? x+12-20=14+8 x= ___ @cj49

OpenStudy (cj49):

x-8=22 x=30

OpenStudy (bossimbacon):

@cj49 thanks can you help me with a couple more?

OpenStudy (bossimbacon):

What is the solution to the equation? 4x-26-3x=16-20 x=___

OpenStudy (cj49):

x-26=-4 x=-4+26 x=22

OpenStudy (bossimbacon):

What is the solution to the equation? 5(x+4)-4x=18-23 x=____

OpenStudy (cj49):

5x+20-4x=-5 x=-20-5 x=-25

OpenStudy (bossimbacon):

What is the solution to the equation? x-7.6+9.4=-8.4 x=___

OpenStudy (bossimbacon):

@cj49

OpenStudy (cj49):

u try this mate ill guide u if u wud be stucked

OpenStudy (bossimbacon):

would the answer be -10.2?

OpenStudy (bossimbacon):

@cj49

OpenStudy (cj49):

correct..